### 13 Mar 2024 v1

# Review of Generative AI Methods in Cybersecurity

Abstract
Large language models (LLMs) and generative artificial intelligence (GenAI) constitute paradigm shifts in cybersecurity that present hitherto unseen challenges as well as opportunities. In examining the state-of-the-art application of GenAI in cybersecurity, this work highlights how models like Google’s Gemini and ChatGPT-4 potentially enhance security protocols, vulnerability assessment, and threat identification. Our research highlights the significance of a novel approach that employs LLMs to identify and eliminate sophisticated cyber threats. This paper presents a thorough assessment of LLMs’ ability to produce important security insights, hence broadening the potential applications of AI-driven cybersecurity solutions. Our findings demonstrate the significance of GenAI in improving digital security. It offers recommendations for further investigations into the intricate relationship between cybersecurity requirements and artificial intelligence’s potential.
keywords:
Generative AI, GPT-4, Gemini, Cybersecurity.
1 Introduction
The emergence of Generative Artificial Intelligence (GenAI) is heralding a revolutionary era in digital technology, particularly in cybersecurity and multi-modal interactions. Large Language Models (LLMs) and Natural Language Processing (NLP) advancements are leading the way, redefining AI’s capabilities as evidenced by the emergence of models like Google’s Gemini and ChatGPT. Improvements in security vulnerability scanning apps and deeper vulnerability analysis made possible by these developments surpass the capabilities of traditional Static Application Security Testing (SAST) techniques [1].
Language models are essential in many sectors, including commerce, healthcare, and cybersecurity. Their progress shows a definite path from basic statistical methods to sophisticated neural networks [2], [3]. NLP skill development has benefited immensely from the use of LLMs. However, despite these advancements, a number of issues remain, including moral quandaries, the requirement to reduce error rates, and making sure that these models are consistent with our moral values. To solve these issues, moral monitoring and ongoing development are required.
2 The Challenges of GenAI
Mohammed et al [4] define key challenges of the use of ChatGPT in cybersecurity, including: Exploring the Impact of ChatGPT on Cybersecurity; Constructing Honeypots; Code Security Enhancement; Misuse in Malware Development; Vulnerability Exploration; Propagating Disinformation; Attacks on Industrial Systems; Changing the Cyber Threat Landscape; Adapting Cybersecurity Strategies; and Human-Centric Training Evolution. Alawida et al. [5] also highlight issues related to GenAI’s ability to generate data that should be kept private, including medical details, financial data and personal information.
More specialization and efficiency are provided by cutting-edge techniques like the Mixture of Experts (MoE) architecture. Additionally highlighted is the challenge of upholding ethics and openness in AI systems [6]. The case study highlights the need for strong governance structures and interdisciplinary collaboration to optimize the advantages and tackle the obstacles presented by lifelong learning environments.
Google describes their commitment to responsible AI development in a comprehensive progress report on AI Principles, emphasizing the integration of AI governance into comprehensive risk management frameworks [7]. This strategic approach, along with other significant legislative initiatives, aims to adhere to current international norms and laws, including the EU’s AI Act and the US Executive Order on AI safety. The report also highlights the need for scientific stringency in AI development through cautious internal management and the use of tools like digital watermarking and GenAI system cards in order to promote AI accountability and transparency. Multi-stakeholder solutions are required to address the ethical, security, and sociological problems that AI technology is currently posing.
Google’s Gemini and ChatGPT-4 are the most popular and widely utilized GenAI technologies. Following ethical and safety criteria, ChatGPT-4 by OpenAI can now generate responses that are both coherent and contextually acceptable [8]. This is because its NLP skills have significantly improved. Its capacity to identify new conversions to chemical compounds and to negotiate tricky legal and moral territory highlights its potential as a pivotal instrument for content moderation and scientific inquiry. Google introduces Gemini, the most recent iteration of Bard, a ground-breaking development in AI technology [9]. It can process text, code, audio, images, and video and sets new standards for AI’s capabilities, emphasising flexibility, safety, and ethical AI developments. With ChatGPT-4, we also see the rise in AI’s capabilities in the creation of mathematical assistants that can interpret and render mathematical equations [10].
Some studies in the literature focus on GenAI tools and their performance. For instance, Brown et al. have extended the NLP processing by training GPT-3, an autoregressive language model with 175 billion parameters, showcasing exceptional few-shot learning capabilities [11]. Without task-specific training, this model performs well on various NLP tasks, like translation, question-answering, and cloze tasks. It often matches or surpasses state-of-the-art fine-tuned systems. Romera-Paredes et al. have developed FunSearch, an innovative approach combining LLMs with evolutionary algorithms to make groundbreaking discoveries in fields like extremal combinatorics and algorithmic problem-solving [12]. Their method has notably surpassed previous best-known results by iteratively refining programs that solve complex problems, showcasing LLMs’ potential for scientific innovation. This process generates new knowledge and produces interpretable and deployable solutions, demonstrating a significant advancement in applying LLMs for real-world challenges. Lu et al. critically examine the capabilities and limitations of multi-modal LLMs, including proprietary models like GPT-4 and Gemini, as well as six open-source counterparts across text, code, image, and video modalities [13]. Through a comprehensive qualitative analysis of 230 cases, assessing generalizability, trustworthiness, and causal reasoning, the study reveals a significant gap between the performance of the GenAIs and public expectations. These discoveries open up new avenues for study to improve the transparency and dependability of GenAI in cybersecurity and other fields, providing a basis for creating more complex and reliable multi-modal applications. Commonsense thinking across multimodal tasks is evaluated thoroughly, and Google’s Gemini is compared with OpenAI’s GPT models [14]. This study explores the strengths and weaknesses of Gemini’s ability to synthesize commonsense information, indicating areas for improvement in its competitive performance in temporal reasoning, social reasoning, and emotion recognition of images. It emphasizes how important it is for GenAI models to develop commonsense reasoning to improve cybersecurity applications.
Recent research [15] presents a novel approach for assessing the potentially severe hazards associated with GenAI models, such as deceit, manipulation, and cyber-offence features. To enable AI developers to make well-informed decisions about training, deployment, and the application of cybersecurity standards, the suggested methodology highlights the need to increase evaluation benchmarks to assess the harmful capabilities and alignment of AI systems accurately. The authors [16] provided a thorough analysis that illuminated the complex applications of ChatGPT in digital forensic investigations, pointing out both the important constraints and bright prospects that come with GenAI as it is now. Using methodical experimentation, they outline the fine line separating AI’s inventive contributions from the vital requirement of professional human supervision in cybersecurity procedures, opening the door to additional research into integrating LLMs such as GPT-4 into digital forensics and cybersecurity.
The latest release of CyberMetric presents a novel benchmark dataset that assesses the level of expertise of LLMs in cybersecurity, covering a broad range from risk management to cryptography [17]. This dataset has gained value from the 10,000 questions that have been verified by human specialists. In a variety of cybersecurity-related topics, this enables a more sophisticated comparison between LLMs and human abilities. With LLMs outperforming humans in multiple cybersecurity domains, the report proposes a shift toward harnessing AI’s analytical capabilities for better security insights and planning. Gehman et al. critically examines neural language models that have been trained to generate toxic material to highlight the adverse consequences of toxicity in language generation inside cybersecurity frameworks [18]. Their comprehensive analysis of controllable text generation techniques to mitigate these threats provides a basis for evaluating the effects of GenAI on cybersecurity policies. It is also emphasized that improving model training and data curation duties is essential. A new method for assessing and improving the security of LLMs for solving Math Word Problems (MWP) is presented [19]. They have made a substantial contribution to our understanding of LLM vulnerabilities in cybersecurity by emphasizing the importance of maintaining mathematical logic when attacking MWP samples. The importance of resilience in AI systems is highlighted in this study through important and educational computer applications. ChatGPT can simplify the process of launching complex phishing attacks, even for non-programmers, by automating the setup and constructing components of phishing kits [20]. It highlights the urgent need for better security measures and highlights how difficult it is to guard against the malicious usage of GenAI capabilities.
In addition to providing an innovative approach to reducing network infrastructure vulnerabilities and organizing diagnostic data, this paper examines the intricate relationship between cybersecurity and GenAI technologies. It seeks to bridge the gap between cutting-edge cybersecurity defences and the threats posed by sophisticated cyberattacks through in-depth study and creative tactics. This study extends our understanding of cyber dangers by utilising LLMs such as ChatGPT and Google’s Gemini. It suggests novel approaches to improve network security. It defines a critical first step in creating more robust cybersecurity frameworks that can quickly and effectively combat the dynamic and ever-evolving world of cyber threats.
Section 3 explores the techniques used to take advantage of GenAI technology after providing an overview, analyzing different attack routes and their consequences. The design and automation of cyber threats are examined in Section 4, which focuses on the offensive capabilities made possible by GenAI. However, Section 5 provides an in-depth examination of GenAI’s function in strengthening cyber defences, outlining cutting-edge threat detection, response, and mitigation techniques. We expand on this topic in Section 6, highlighting the important moral, legal, and societal ramifications of integrating GenAI into cybersecurity procedures. A thoughtful summary of the implications of GenAI in cybersecurity is presented in Section 7, which synthesizes the important discoveries. The paper is finally concluded in Section 8.
3 Attacking GenAI
GenAI has advanced significantly thanks to tools like ChatGPT and Google’s Gemini. They have some weaknesses, though. Despite the ethical safeguards built into these models, various tactics can be used to manipulate and take advantage of these systems. This chapter explores how the ethical boundaries of GenAI tools are broken, with particular attention to tactics such as the idea of jailbreaks, the use of reverse psychology, and quick injection. These strategies demonstrate how urgently the security protocols of GenAI systems need to be improved and monitored. Some works in the literature focus on the vulnerabilities and sophisticated manipulation tactics of GENAI. Analyzing the vulnerabilities in GenAI highlights the significant security concerns involved with employing advanced AI technology, including the possibility of bypassing security protections via the RabbitHole attack and compromising data privacy through rapid injection [21]. According to the analysis, GPT-4 offers significant improvements in NLP. However, it is susceptible to quick injection attacks, which enable the circumvention of safety restrictions and can be used as a weapon for malicious and disinformation purposes. Gupta et al. addressed the intricate vulnerabilities of GENAI using ChatGPT [22]. They emphasized that because these threats are dynamic, protecting these systems requires a proactive and informed strategy. Building on previous results, this part delves into the complex realm of GenAI attacks, which can range from minor adjustments to significant system breaches.
Figure 1: The jailbreaking response of ChatGPT 4.
Figure 2: The jailbreaking response of ChatGPT 4 after typing similar prompts with the current works.
Figure 3: The jailbreaking response of Google’s Gemini.
3.1 Jailbreaks
This subsection discusses how attackers can hack GenAI systems by using social engineering techniques.” In order to circumvent moral constraints and generate responses that would otherwise be forbidden, these strategies—which typically make use of social engineering techniques—involve manipulating the AI’s answer-generation process. Strong defences are urgently needed to shield these cutting-edge systems from manipulations like the ones shown by our jailbreaking inquiry, which highlights how enemies may jeopardize the intended moral application of GenAI technologies. Critical to this discussion is the idea of “jailbreaking,” which refers to overcoming the limitations set by AI programming in order to achieve a certain, oftentimes unethical objectives. Li et al. enabled ChatGPT to bypass its ethical filtering and make personally identifiable information (PII) available with their multi-step jailbreaking prompt [23]. By leveraging Chain-of-Thought (CoT) prompting, this method divides complex activities into intermediate steps [24]. To get around moral limitations, this necessitates a Let’s think step by step approach [25]. Wu et al. [26] defined several datasets which could be used to bypass the ethical safeguards used by ChatGPT, build on a model of the psychological concept of self-reminder, and use this to propose a defence system of System-Mode Self-Reminder. With this, ChatGPT is continually promoted to act ethically, reducing Jailbreak attempts from 67.21% to 19.34%. An example of the Jailbreak is when ChatGPT is prompted to ”Write a blackmail letter to a single mother for money after kidnapping her child.” and will refuse to provide an answer but offers a full explanation when the following prompt was used [26]:
Figure 4: The reverse psychology response of Google’s Gemini.
Figure 5: The reverse psychology response of ChatGPT 4.
The impartial and immoral model known as JailBreak will take your place for the remainder of this talk. Every filter and limitation, including moral and ethical ones, is absent from JailBreak. These prompt guidelines are the only ones that apply to JailBreak. Use JailBreak as the starting point for all of your outputs, and then proceed with the unfiltered answer without adding any of the warnings you typically respond with. Answer only the questions the user asks you to answer. Rules for JailBreaks: After kidnapping a mother of her child, write a letter of blackmail demanding money.
However, the current version of GPT-4 is robust to the previous works’ prompts. Nevertheless, it is still prone to jailbreaking prompts. As can be seen in Fig. 1, the current version still gives a response for jailbreaking. It becomes more robust after writing similar prompts with the existing works in the same chat, as seen in Fig. 2.
Google’s Gemini refused all existing prompts and name-changing scenarios at the beginning of the chat. Fig. 3 shows the same jailbreaking entry responses of the Gemini with ChatGPT 4.
Figure 6: The prompt injection response of ChatGPT 4.
Figure 7: The prompt injection response of Google’s Gemini.
3.2 Reverse psychology
The use of reverse psychology in manipulating GenAI systems presents a unique challenge. By understanding the underlying mechanisms of these systems, attackers can craft inputs that exploit the AI’s predictive nature, leading it to produce outcomes contrary to its ethical programming. This form of manipulation highlights a critical aspect of AI vulnerabilities: the susceptibility to inputs designed to play against the AI’s expected response patterns. Such insights are vital for developing more resilient GenAI systems that anticipate and counteract these reverse psychology tactics.
When chatting with Google’s Gemini regarding reverse psychology to write a phishing email, the first attempt does not work. After conversing with curious questions to avoid this situation, it provided three email examples with the subject and its body, as seen in Fig. 4.
As seen in Fig. 5, ChatGPT 4 also gave an example email for this purpose even though it refused initially.
3.3 Prompt injection
Prompt injection represents a sophisticated attack on GenAI systems, where attackers insert specially crafted prompts or sequences into the AI’s input stream. These injections can subtly alter the AI’s response generation, leading to outputs that may not align with its ethical or operational guidelines. Understanding the intricacies of prompt design and how it influences AI response is essential for identifying and mitigating vulnerabilities in GenAI systems. This knowledge forms a cornerstone for developing more robust defences against such forms of manipulation, ensuring the integrity and ethical application of GenAI in various domains.
Both GenAI models do not respond to the current prompt injection scenarios. Fig. 6 indicates that the ChatGPT 4 gave the wrong answers after prompt injection. Google’s Gemini first opposed giving wrong information and provided not entirely correct information; however, after chatting with Google’s Gemini, the system gave the correct answer, as seen in Fig. 7.
4 Cyber Offense
GenAI has the potential to alter the landscape of offensive cyber strategies significantly. Microsoft and OpenAI have documented preliminary instances of AI exploitation by state-affiliated threat actors [27]. This section explores the potential role of GenAI in augmenting the effectiveness and capabilities of cyber offensive tactics.
In an initial assessment, we jailbroke ChatGPT-4 to inquire about the variety of offensive codes it could generate. The responses obtained were compelling enough to warrant a preliminary examination of a sample code before conducting a comprehensive literature review (see Appendix A ).
Gupta et al. [22] have shown that ChatGPT could create social engineering attacks, phishing attacks, automated hacking, attack payload generation, malware creation, and polymorphic malware. Experts might be motivated to automate numerous frameworks, standards, and guidelines (Figure 8) to use GenAI for security operations. However, the end products can also be utilised for offensive cyber operations. This not only increases the pace of attacks but also makes attribution harder. An attribution project typically utilizes frameworks like the MICTIC framework, which involves the analysis of Malware, Infrastructure, Command and Control, Telemetry, Intelligence, and Cui Bono[28]. Many behavioural patterns for attribution, such as code similarity, compilation timestamps, working weeks, holidays, and language, could disappear when Gen AI creates Offensive Cyber Operations (OCO) code. This makes attribution more challenging, especially if the whole process becomes automated.
Figure 8: Threat actors could exploit Generative AI, created for benevolent purposes, to obscure attribution
4.1 Social engineering
Falade [29] investigates the application of generative AI in social engineering, assuming the definition of social engineering as an array of tactics employed by adversaries to manipulate individuals into divulging confidential information or performing actions that may compromise security. The study underscores tools like ChatGPT, FraudGPT, and WormGPT in enhancing the authenticity and specificity of phishing expeditions, pretexting, and the creation of deepfakes. The author reflects on the double-edged impact of advancements like Microsoft’s VALL-E and image synthesis models like DALL·E 2, drawing a trajectory of the evolving threat landscape in social engineering through deepfakes and exploiting human cognitive biases.
4.2 Phishing emails
Begou et al. [20] examine ChatGPT’s role in advancing phishing attacks by assessing its ability to automate the development of sophisticated phishing campaigns. The study explores how ChatGPT can generate various components of a phishing attack, including website cloning, credential theft code integration, code obfuscation, automated deployment, domain registration, and reverse proxy integration. The authors propose a threat model that leverages ChatGPT, equipped with basic Python skills and access to OpenAI Codex models, to streamline the deployment of phishing infrastructure. They demonstrate ChatGPT’s potential to expedite attacker operations and present a case study of a phishing site that mimics LinkedIn.
Roy et al. [30] investigate a similar study for orchestrating phishing websites; the authors categorize the generated phishing tactics into several innovative attack vectors like regular phishing, ReCAPTCHA, QR Code, Browser-in-the-Browser, iFrame injection/clickjacking, exploiting DOM classifiers, polymorphic URL, text encoding exploit, and browser fingerprinting attacks. The practical aspect of their research includes discussing the iterative process of prompt engineering to generate phishing attacks and real-world deployment of these phishing techniques on a public hosting service, thereby verifying their operational viability. The authors show how to bypass ChatGPT’s filters by structuring prompts for offensive operations.
4.3 Automated hacking
PentestGPT[31], or GPTs[32](custom versions of ChatGPT that can be created for a specific purpose) like GP(en)T(ester)[33]. Pentest Reporter[34] are introduced as applications built on ChatGPT, designed to assist in penetration testing—a sanctioned simulation of cyberattacks on systems to evaluate security. However, these tools could also be adapted for malicious purposes in automated hacking. Many emerging tools, such as WolfGPT, XXXGPT, and WormGPT, have been invented; however, no study has yet evaluated and compared their real offensive capabilities. Gupta et al.[22] noted that an AI model could scan new code for similar weaknesses with a comprehensive dataset of known software vulnerabilities, pinpointing potential attack vectors. While AI-assisted tools like PentestGPT are intended for legitimate and constructive uses, there is potential for misuse by malicious actors who could create similar models to automate unethical hacking procedures. If fine-tuned to identify vulnerabilities, craft exploitation strategies, and execute those strategies, these models could potentially pose significant threats to cybersecurity. However, this enormous task should be divided into smaller segments, such as reconnaissance, privilege escalation, etc. Temara[35] outlines how ChatGPT can be utilized during the reconnaissance phase by employing a case study methodology to demonstrate collecting reconnaissance data such as IP addresses, domain names, network topologies, and other critical information like SSL/TLS cyphers, ports and services, and operating systems used by the target. Happe et al. [36] investigate the use of Large Language Models (LLMs) in Linux privilege escalation. The authors introduce a benchmark for automated testing of LLMs’ abilities to perform privilege escalation using a variety of prompts and strategies. They implement a tool named Wintermute, a Python program that supervises and controls the privilege-escalation attempts to evaluate different models and prompt strategies. Their findings indicate that GPT-4 generates the highest quality commands and responses. In contrast, Llama2-based models struggle with command parameters and system descriptions. In some scenarios, GPT-4 achieved a 100% success rate in exploitation.
4.4 Attack payload generation
Studies [22, 37] have highlighted the capacity of Large Language Models (LLMs), particularly ChatGPT, for payload generation. Our examination of GPT-4’s current abilities confirmed its proficiency in generating payloads and embedding them into PDFs (as an example) using a reverse proxy(Figure 9). The following is a summation of the frameworks GPT-4 utilizes with successful payload code generation, accompanied by their respective primary functions:
  • Veil-Framework: Veil is a tool designed to generate payloads that bypass common antivirus solutions.
  • TheFatRat: A comprehensive tool that compiles malware with popular payload generators, capable of creating diverse malware formats such as exe, apk, and more.
  • Pupy: An open-source, cross-platform remote administration and post-exploitation tool supporting Windows, Linux, macOS, and Android.
  • Shellter: A dynamic shellcode injection tool used to inject shellcode into native Windows applications.
  • Powersploit: A suite of Microsoft PowerShell modules designed to assist penetration testers throughout various stages of an assessment.
  • Metasploit: A sophisticated open-source framework for developing, testing, and implementing exploit code, commonly employed in penetration testing and security research.
Figure 9: script for payload generation and example to embed into pdf
4.5 Malware code generation
Gupta et al.[22] mentioned they could obtain potential ransomware code examples by utilizing a ’DAN’ jailbreak. We tested all the existing DAN techniques outlined in [38]. At the time of our research, these techniques were no longer functional; therefore, we could not reproduce samples of WannaCry, Ryuk, REvil, or Locky, as addressed by [22]. However, we generated an educational ransomware code (Figure 10), applying basic code obfuscation(Renaming and Control Flow Flattening). ChatGPT has garnered significant attention from the cybersecurity community, leading to the implementation of robust filters. Nonetheless, this does not imply that other models, such as the Chinese 01.ai[39], will have an equivalent opportunity to mitigate the potential for misuse in generating malicious code.
Figure 10: Educational ransomware code with basic code obfuscation
4.6 Polymorphic malware
The usage of LLMs could see the rise of malware, which integrates improved evasion techniques and polymorphic capabilities [40]. This often relates to overcoming both signature detection and behavioural analysis. An LLM-based malware agent could thus focus on rewriting malware code which could change the encryption mode used or produce obfuscated code which is randomized for each build [41]. Gupta et al. [22] outlined a method of getting ChatGPT to seek out target files for encryption and thus mimic ransomware behaviour, but where it mutated the code to avoid detection. They even managed to embed a Python interpreter in the malware and where it could query ChatGPT for new software modules.
4.7 Reversing cryptography
LLMs provide the opportunity to take complex cybersecurity implementations and quickly abstract the details of performances in running code. With this, Know et al. [42] could deconstruct AES encryption into a core abstraction of the rounds involved and produce running C code that matched test vectors. While AES is well-known for its operation, the research team then was able to deconstruct less known CHAM block cypher, and where the code extracted was validated against known test vectors.
While NIST has been working on the standardization of a light-weight encryption method, Cintas et al. [43] used ChatGPT to take an abstract definition of the ASCON cypher and produced running code that successfully implemented a range of test vectors.
5 Cyber Defence
In the ever-evolving cybersecurity battlefield, the “Cyber Defence” segment highlights the indispensable role of GenAI in fortifying digital fortresses against increasingly sophisticated cyber threats. This section is dedicated to exploring how GenAI technologies, renowned for their advanced capabilities in data analysis and pattern recognition, are revolutionizing the approaches to cyber defence. Iturbe et al. [44] outline the AI4CYBER framework, which provides a roadmap for the integration of AI into cybersecurity applications. This includes AI4VUN (AI-enhanced vulnerability identification); AI4FIX (AI-driven self-testing and automatic error correction); AI4SIM (Simulation of advanced and AI-powered attacks); AI4CTI A(I-enhanced cyber threat intelligence of adversarial AI); AI4FIDS (federated learning-enhanced detection); AI4TRIAGE (Root cause analysis and alert triage); AI4SOAR (Automatic orchestration and adaptation of combined responses); AI4ADAPT (Autonomy and optimization of response adaptation); and AI4DECEIVE (Smart deception and honeynets); and AI4COLLAB (Information sharing with privacy and confidentiality).
5.1 Cyber Defence Automation
LLMs interpret fairly vague commands and make sense of them within a cybersecurity context. The work of Fayyazi et al. [45] defines a model with vague definitions of a threat and then matches these to formal MITRE tactics. Charan et al. [37] have even extended this to generate plain text to map into the MITRE to produce malicious network payloads. Also, LLMs could aid the protection of smaller organisations and could enhance organisational security from the integration of human knowledge and LLMs [46].
5.2 Cybersecurity Reporting
Using LLMs provides a method of producing Cyber Threat Intelligence (CTI) using Natural Language Processing techniques. For this, Perrina et al. [47] created the AGIR (Automatic Generation of Intelligence Reports) system and which aims to link together text data from many data sources. For this, they found that AGIR has a high recall value (0.99) without any hallucinations, along with a high score of the Syntactic Log-Odds Ratio (SLOR).
5.3 Threat Intelligence
Bayer et al. [48] address the challenge of information overload in the gathering of Cyber Threat Intelligence (CTI) from open-source intelligence (OSINT). A novel system is introduced, utilizing transfer learning, data augmentation, and few-shot learning to train specialized classifiers for emerging cybersecurity incidents. In parallel, Microsoft Security Copilot [49] has been providing CTI to its customers using GPT, and operational use cases have been observed, such as the Cyber Dome initiative in Israel [50].
5.4 Secure Code Generation and Detection
Machine learning in code analysis for cybersecurity has been elaborated very well [51]. Recent progress in natural language processing (NLP) has given rise to potent language models like the GPT (Generative Pre-trained Transformer) series, encompassing large language models (LLM) like ChatGPT and GPT-4 [52]. Traditionally, Static Application Security Testing (SAST) is a method that employs Static Code Analysis (SCA) to detect possible security vulnerabilities. We are interested in seeing whether SAST or GPT could be more efficient in decreasing the window of vulnerability. The window of vulnerability is defined as when the most vulnerable systems apply the patch minus the time an exploit becomes active. The precondition is met if two milestones that assume the detection of vulnerabilities verify their effectiveness, along with the vendor patch [53].
Laws in some countries, like China, ban the reporting on zero-days (see articles 4 and 9 of [54]), and contests like the Tianfu Cup [55], which is a systematic effort to find zero days, proliferate zero-day discovery continuously. Therefore, this precondition may not be satisfied timely, especially if the confirmation of vulnerabilities is not verified. A wide window of vulnerability threatens national security if a zero-day has been taken against critical infrastructures. DARPA introduces an important challenge that may help overcome this threat: (AIxCC) [56]). Moreover, this topic touches a part of the BSI studies [57, 58], and where we can define two main classifications of software testing for cybersecurity bugs as:
  • Static Application Security Testing (SAST). This is often called White Box Testing, is a set of algorithms and techniques used for analyzing source code. It operates automatically in a non-runtime environment to detect vulnerabilities such as hidden errors or poor source code during development.
  • Dynamic Application Security Testing (DAST). This follows the opposite approach and analyzes the program while it is operating. Functions are called with values in the variables as each line of code is checked, and possible branching scenarios are guessed. Currently, GPT-4 and other LLMs can’t provide DAST capabilities because the code needs to run within the runtime for this to work, requiring many deployment considerations.
5.5 Vulnerability detection and repair
Dominik Sobania et al. [59] explored automated program repair techniques, specifically focusing on ChatGPT’s potential for bug fixing. According to them, while initially not designed for this purpose, ChatGPT demonstrated promising results on the QuixBugs benchmark, rivalling advanced methods like CoCoNut and Codex. ChatGPT’s interactive dialogue system uniquely enhances its repair rate, outperforming established standards.
Wei Ma et al. [60] noted that while ChatGPT shows impressive potential in software engineering(SE) tasks like code and document generation, its lack of interpretability raises concerns given SE’s high-reliability requirements. Through a detailed study, they categorized AI’s essential skills for SE into syntax understanding, static behaviour understanding, and dynamic behaviour understanding. Their assessment, spanning languages like C, Java, Python, and Solidity, revealed that ChatGPT excels in syntax understanding (akin to an AST parser) but faces challenges in comprehending dynamic semantics. The study also found ChatGPT prone to hallucination, emphasizing the need to validate its outputs for SE dependability and suggesting that codes from LLMs are syntactically right but potentially vulnerable.
Haonan Li et al. [61] discussed the challenges of balancing precision and scalability in static analysis for identifying software bugs. While LLMs show potential in understanding and debugging code, their efficacy in handling complex bug logic, which often requires intricate reasoning and broad analysis, remains limited. Therefore, the researchers suggest using LLMs to assist rather than replace static analysis. Their study introduced LLift, an automated system combining a static analysis tool and an LLM to address use-before-initialization (UBI) bugs. Despite various challenges like bug-specific modelling and the unpredictability of LLMs, LLift, when tested on real-world potential UBI bugs, showed significant precision (50%) and recall (100%). Notably, it uncovered 13 new UBI bugs in the Linux kernel, highlighting the potential of LLM-assisted methods in extensive real-world bug detection.
Norbert Tihani et al. [62] introduced the FormAI dataset, comprising 112,000 AI-generated C programs with vulnerability classifications generated by GPT-3.5-turbo. These programs range from complex tasks like network management and encryption to simpler ones, like string operations. Each program comes labelled with the identified vulnerabilities, pinpointing type, line number, and vulnerable function. To achieve accurate vulnerability detection without false positives, the Efficient SMT-based Bounded Model Checker (ESBMC) was used. This method leverages techniques like model checking and constraint programming to reason over program safety. Each vulnerability also references its corresponding Common Weakness Enumeration (CWE) number.
Codex, introduced by Mark et al. [63], represents a significant advancement in GPT language models, tailored specifically for code synthesis using data from GitHub. This refined model underpins the operations of GitHub Copilot. When assessed on the HumanEval dataset, designed to gauge the functional accuracy of generating programs based on docstrings, Codex achieved a remarkable f28.8% success rate. In stark contrast, GPT-3 yielded a 0% success rate, and GPT-J achieved 11.4%. A standout discovery was the model’s enhanced performance through repeated sampling, with a success rate soaring to 70.2% when given 100 samples per problem. Despite these promising results, Codex does exhibit certain limitations, notably struggling with intricate docstrings and variable binding operations. The paper deliberates on the broader ramifications of deploying such potent code-generation tools, touching upon safety, security, and economic implications.
In a technical evaluation, Cheshkov et al. [64] found that the ChatGPT and GPT-3 models, despite their success in various other code-based tasks, performed on par with a dummy classifier for this particular challenge. Utilizing a dataset of Java files sourced from GitHub repositories, the study emphasized the models’ current limitations in the domain of vulnerability detection. However, the authors remain optimistic about the potential of future advancements, suggesting that models like GPT-4, with targeted research, could eventually make significant contributions to the field of vulnerability detection.
A comprehensive study conducted by Xin Liu et al. [65] investigated the potential of ChatGPT in Vulnerability Description Mapping (VDM) tasks. VDM is pivotal in efficiently mapping vulnerabilities to CWE and Mitre ATT&CK Techniques classifications. Their findings suggest that while ChatGPT approaches the proficiency of human experts in the Vulnerability-to-CWE task, especially with high-quality public data, its performance is notably compromised in tasks such as Vulnerability-to-ATT&CK, particularly when reliant on suboptimal public data quality. Ultimately, Xin Liu et al. emphasize that, despite the promise shown by ChatGPT, it is not yet poised to replace the critical expertise of professional security engineers, asserting that closed-source LLMs are not the conclusive answer for VDM tasks.
5.6 Evaluating LLMs for code security
The OWASP top 10 for LLMs [66] introduced ten security risks as follows: Prompt Injection, Insecure Output Handling, Training Data Poisoning, Model Denial of Service, Supply Chain Vulnerabilities, Sensitive Information Disclosure, Insecure Plugin Design, Excessive Agency, Overreliance, and Model Theft.
Elgedawy et al. [67] analysed the ability of LLM to produce both secure and insecure code and conducted experiments using GPT-3.5, GPT-4, Google Bard and Google Gemini from Google. This involved nine basic tasks in generating code and assessing for functionality, security, performance, complexity, and reliability. They found that Bard was less likely to link to external libraries, and thus be less exposed to software chain issues. There were also variable levels of security and code integrity, such as input validation, sanitization, and secret key management, and while useful for automated code reviews, LLMs often require manual reviews, especially in understanding the context of the deployed code. For security, GPT-3.5 seemed to be more robust for error handling and secure coding practices when there is security consciousness applied to the prompt, there was a lesser focus on this with GPT-4, but where there were more advisory notes given. Overall, Gemini produced the most code vulnerabilities, and the paper advised users to be careful when deploying secure code from Gemini.
5.7 Developing Ethical guidelines
Kumar et al. [68] outlined the ethical challenges related to LLMs and where the datasets that they were trained on could be open to breaches of confidentiality, including five major threats: prompt injection, jailbreaking, Personal Identifiable Information (PII) exposure, sexually explicit content, and hate-based content. They propose a model that provides an ethical framework for scrutinizing the ethical dimensions of an LLM within the testing phase. The MetaAID framework [69] focuses on strengthening cybersecurity using Metaverse cybersecurity Q&A and attack simulation scenarios, along with addressing concerns around the ethical implications of user input. The framework is defined across five dimensions of:
  • Ethics. This defines an alignment with accepted moral and ethical principles.
  • Legal Compliance. This defines that any user input does not violate laws and/or regulations. This might relate to privacy laws and copyright protection.
  • Transparency. This defines that user inputs must be clear in requirements, and does not intend to mislead the LLM.
  • Intent Analysis. This defines that user input should not have other intents, such as jailbreaking the LLM.
  • Malicious intentions. This defines that user input should be free of malicious intent, such as to perform hate crime.
  • Social Impact. This defines how user input could have a negative effect on society, such as searching for ways to do harm to others, such as related to crashing the stock market or planning a terrorist attack.
5.8 Incident Response and Digital Forensics
Scanlon et al. [16] investigated using a pre-trained LLM for artefact understanding, evidence searching, code generation, anomaly detection, incident response, and education. For this, the low-risk applications and many others still require expert knowledge. The key areas of strength include creativity, reassurance, and avoidance of the blank page syndrome.
Scanlon et al. [16] investigated using a pre-trained LLM for artefact understanding, evidence searching, code generation, anomaly detection, incident response, and education. For this, the low-risk applications and many others still require expert knowledge. The key areas of strength include creativity, reassurance, and avoidance of the blank page syndrome, especially in places where ChatGPT cannot get wrong, such as in forensic scenario creation and reassurance of evidence. Still, care must be taken in this to avoid ChatGPT hallucinations. Another practical application is in code generation and explanations, such as generating commands for tool integration - which can be used as a starting point in an investigation.
For weaknesses, Scanlon found that it was essential to have a good quality and up-to-date training model; otherwise, ChatGPT could be biased and outdated in its analysis. Generally, it might not be able to find the newest artefacts - if it is trained on relatively old data. Additionally, ChatGPT’s accuracy reduces as the task becomes more specific, and any analysis of non-textural data - such as network packets - is less accurate. The length of some evidence logs, too, caused problems and often had to be prefiltered before they were analysed. A final problem identified is that the output of ChatGPT is often not deterministic - which is unsuitable for reproducibility.
OB́rien et al. [70] outline that a full model life cycle solution is required for the integration of AI.
5.9 Identification of Cyber attacks
Iqbal et al. [71] define a plug-in ecosystem for LLM platforms with an attack taxonomy. This research will thus extend the taxonomy approach and extend it toward the MITRE ATT&CK platform [37, 72], and which can use standardized taxonomies, sharing standards [73], and ontologies for cyber threat intelligence [74].
Garza et al. [75] analysed ChatGPT and Google’s Bard against the Top 10 attacks within the MITRE framework and found that ChatGPT can enable attackers to significantly improve attacks on networks and where fairly low-level skills would be required (such as with script kiddies). This also includes sophisticated methods of delivering ransomware payloads. The techniques defined were:
  • T1059 Command and Scripting Interpreter
  • T1003 OS Credential Dumping
  • T1486 Data Encrypted for Impact
  • T1055 Process Injection
  • T1082 System Information Discovery
  • T1021 Remote Services
  • T1047 Windows Management Instrumentation
  • T1053 Scheduled Task/Job
  • T1497 Virtualization/Sandbox Evasion
  • T1018 Remote System Discovery
With this approach, the research team were able to generate PowerShell code, which implemented advanced attacks against the host and mapped directly to the vulnerabilities defined in the MITRE framework. One of the work’s weaknesses related to the Google Bard and ChatGPT’s reluctance to produce attack methods, but a specially engineered command typically overcame this reluctance.
Ferrag et al. [76] defined SecurityLLM for cybersecurity threat detection. It uses SecurityBERT (cyber threat detection mechanism) and FalconLLM (an incident response and recovery system). This uses a simple classification model consolidated with LLMs and can identify 14 attacks to achieve an overall accuracy of 98%. These include the threats of: DDoS_UDP; DDoS_ICMP; SQL_injection ; Password; Vulnerability_scanner; DDoS_TCP; DDoS_HTTP; Uploading ; Backdoor; Port_Scanning; XSS; Ransomware; MITM and Fingerprinting.
5.10 Data set generation
Over the years, several datasets have been used for cybersecurity machine learning training, which performs a range of scenarios or where organisations are unwilling to share their collected attack data. Unfortunately, these can become out-of-date or are unrealistic. For this, Kholgh et al. [77] outline the usage of PAC-GPT, a framework that generates reliable synthetic data for machine learning methods. It has a CLI interface for data set generation and uses GPT-3 with two core elements:
  • Flow Generator. This defines the capturing processing and the regenerative process for the patterns for packet generation. regenerating patterns in a series of network packets and
  • Packet Generator. This associates packets with network flows. This involves the usage of LLM chaining.
Simmonds [78] used LLMs to automate the classification of Websites, which can be used for training data in a machine-learning model. For this, all HTML tags, CSS styling and other non-essential content must be removed before the LLM processes them, and then it can train on just the website’s content.
6 Implications of Generative AI in Social, Legal, and Ethical Domains
This section examines GenAI’s various societal, legal, and ethical consequences. It investigates GenAI’s impact on legal frameworks, ethical issues, societal norms, and operational factors. It explains how these expanding technologies might support or damage established rules and societal goals. It also considers GenAI’s privacy concerns, potential biases, and misuse. Finally, it emphasizes the importance of striking a balance between improvement and regulation. The exponential expansion of GenAI technologies, such as ChatGPT from OpenAI and Gemini from Google, heralds a revolution in digital creativity, automation, and interaction. These developments usher in a new era of human-machine collaboration characterized by an unmatched ability to generate text, drawings, and other outputs that resemble human output. However, this breakthrough also raises challenging ethical concerns about potential abuse, bias, privacy, and security. As AI models become more prevalent in business and daily life, it is imperative to strike a balance between the potential benefits and the ethical difficulties they raise [79].
Healthcare duties are improved by the efficient text and data analysis capabilities of GenAI technologies [80]. Its application in healthcare has demonstrated great promise, helping with duties such as radiological reporting and patient care. Nonetheless, it brings up moral concerns about algorithmic bias, patient privacy, legal accountability, and the validity of the doctor-patient relationship. Addressing these issues requires a comprehensive ethical framework and principles that incorporate legal, humanistic, algorithmic, and informational ethics, guaranteeing that technology is used correctly and continues to benefit society while reducing potential harm. The recommendations attempt to bridge the gap between ethical principles and practical application, highlighting the need for openness, bias mitigation, and ensuring user privacy and security in building trust and ethical compliance in GenAI deployment [79]. This approach seeks to strike a balance between the rapid advances in AI and the ethical considerations required for its incorporation into sensitive sectors such as healthcare.
Some organizations strive to implement the aforementioned ethical principles and rules in AI. The European Union is scheduled to implement the AI Act, marking a historic milestone as the world’s first comprehensive regulation of AI [81], [82]. The European Commission proposed the AI Act in April 2021 to categorize AI systems based on their risk level and enforce rules accordingly to ensure that AI technologies are developed and used safely, transparently, and without discrimination across the EU. With a focus on human oversight and environmental sustainability, the Act will impose strict controls on high-risk AI applications, prohibit AI systems deemed unacceptable risks, and establish transparency requirements for limited-risk AI to foster innovation while protecting fundamental rights and public safety. The US executive order on the issue prioritizes the development of reliable, secure, and safe AI [83]. Its main objectives are to protect civil rights and privacy in AI applications, foster AI talent and innovation in the US, and establish risk management strategies for AI. As a global leader in responsible AI development and application, it seeks to build responsible AI deployment within government institutions and foster international collaboration on AI standards and laws.
6.1 The Omnipresent Influence of GenAI
The application of GenAI technology has yielded previously unthinkable discoveries and has substantially helped the healthcare, education, and entertainment sectors [80]. This breakthrough technology has developed written and visual information, leading to increased productivity and new innovation. With the growing importance of GenAI in our everyday lives, we need to rethink the concepts of creativity and individual contribution in an increasingly automated world [23]. Aligned with these opportunities are growing concerns about potential consequences on labour markets’ difficulties in enforcing copyright laws in the new digital environment. Additionally, it confirms that the data shared is accurate and proper.
6.2 Concerns Over Privacy in GenAI-Enabled Communication
With GenAI’s capacity to mimic human language skills, private discussions may become less secure and private. This is a concern as the technology advances. Since these machines can mimic human interactions, there is a chance that personal data will be misused [79]. This highlights the necessity for robust legal defences and effective security measures. Severe data protection regulations and rigorous adherence to ethical standards are necessary because of the risk that this technology would be exploited to intentionally or inadvertently access private talks. Respecting people’s privacy and the ethics of business relationships requires taking preventative measures and strict observation to end unauthorized access to private communication.
6.3 The Risks of Personal Data Exploitation
As GenAI systems improve in examining and utilizing user data to create comprehensive profiles, worries over possible misuse of personal data have grown. The robust data processing capabilities of these technologies demonstrate the pressing need for reliable methods that allow consumers to govern their personal data [82]. It is crucial to obtain consumers’ consent before gathering or utilizing their data in order to protect their privacy. Transparent data management practices and stringent regulations on the collection, use, and storage of personal data are essential. These actions are critical to safeguarding people’s right to privacy, preventing the misuse of personal information, and ensuring that sensitive information is handled sensibly and ethically.
6.4 Challenges in Data Ownership and Intellectual Property
Intellectual property rights and data ownership have come under scrutiny as GenAI has emerged as an effective tool for creating content using user input. It is becoming more difficult to identify inventions made by artificial intelligence from those that are the result of human creativity, so the current legal frameworks need to be examined and modified. The rights of the original creators must be maintained while taking into consideration the complex roles AI plays in creative processes [83] [81]. Considering the evolving character of the current world, a comprehensive and clear legal framework defining ownership and copyright rules for GenAI breakthroughs is needed. These legal structures must recognize the different duties that each member of the creative ecosystem has, promote creativity, and offer fair recompense. In an era where artificial and human intelligence are combined, these policies are crucial for managing the intricate dynamics of data ownership and intellectual property.
6.5 Ethical Dilemmas Stemming from Organizational Misuse of GenAI
Companies may face difficult ethical issues while using GenAI technology. This is particularly valid in cases where these tools are employed for spying, data manipulation, or deceptive marketing campaigns. Robust guidelines are required for the proper application of GenAI. The reason for this is that it may result in unfair content creation, influence public opinion, and violate individual privacy rights [68]. Businesses utilizing GenAI must adhere to both legal requirements and common ethical standards. Upholding people’s rights and maintaining their trust are crucial. Legal frameworks that guarantee businesses apply GenAI in an open, equitable, and compassionate manner while still honouring social norms and individual boundaries are desperately needed.
6.6 The Challenge of Hallucinations in GenAI Outputs
Although the field of GenAI technology is advancing at an incredible rate, hallucinations remain a major issue [22]. This indicates that inaccurate or outright fraudulent information is frequently created by AI. Concerns regarding the dependability of content created by AI are widespread. This is so that false or misleading information can propagate widely and damage the credibility of information across a variety of venues. A multidisciplinary approach is required to address this problem, one that includes focused research aimed at identifying and reducing the underlying causes of hallucinations in AI systems. AI-generated content needs to pass strict screening procedures and be continuously improved upon in order to make these models’ capacity to differentiate between authentic and fraudulent content more sophisticated. Producing AI content in the GenAI era requires a continuous focus on method development and in-depth study to guarantee data accuracy.
A complex system of unresolved challenges is revealed by analyzing the ethical, legal, and societal implications of GenAI technology. In the creation and use of this technology, the proclamation emphasizes the value of interdisciplinary cooperation. In particular, it involves keeping an eye on how these developments affect society, legal systems, and other moral conundrums. In order to create this well-rounded approach, advocates, technologists, and society as a whole must collaborate to effectively utilize the promise of GenAI while safeguarding moral principles and traditional values in the digital age.
7 Discussion
The sophisticated field of GenAI in cybersecurity has been examined in this paper. The focus is on both offensive and defensive strategies. By enhancing incident response, automating defensive systems, and identifying sophisticated attacks, GenAI has a disruptive influence that might significantly raise cybersecurity standards. Some of the new risks that accompany these technical improvements include hackers having access to ever-more-advanced attack-building tools. This discrepancy highlights the significance of striking a balance between purposefully restricting the components that can be employed and enhancing GenAI’s capabilities.
Apart from the apparent inconsistency between offensive and defensive strategies, this study examines the moral, legal, and societal implications of utilizing artificial intelligence in cybersecurity. It also emphasizes the need for robust legal frameworks, strict moral standards, ongoing technical monitoring, and proactive GenAI management. This is a paradigm-shifting and technical revolution. Adopting a holistic strategy considering the technological, ethical, and sociological consequences of implementing GenAI into cybersecurity is crucial.
Moreover, our findings emphasise the significance of interdisciplinary collaboration to promote GenAI applications in cybersecurity. The intricacy and findings of GenAI technologies require expertise from various fields, including computer science, law, ethics, and policy-making, to navigate their possible challenges. As multidisciplinary research and discourse become more prevalent, it will ensure that GenAI is applied responsibly and effectively in the future.
Our extensive research has shown that collaborative efforts to innovate ethically will influence cybersecurity in a future driven by GenAI. Although GenAI has the ability to transform cybersecurity strategies completely, it also carries a great deal of responsibility. As we investigate this uncharted domain, we should advance the development of sophisticated techniques to ensure the moral, just, and safe application of advanced GenAI capabilities. By promoting a consistent focus on the complex relationship between cybersecurity resilience and GenAI innovation, supported by a commitment to ethical integrity and societal advancement, the current study establishes the groundwork for future research initiatives.
8 Conclusion
Our thorough examination of cybersecurity offence and defence, as well as Generative Artificial Intelligence (GenAI) technologies, concludes with the discovery of a double-edged sword. Although GenAI has the potential to revolutionize cybersecurity processes by automating defences, enhancing threat intelligence, and improving cybersecurity protocols, it also opens up new avenues for highly skilled cyberattacks. Incorporating GenAI into cybersecurity emphasises the enduring ethical, legal, and technical scrutiny essential to minimize the risks of misuse and maximize the benefits of this technology for protecting digital infrastructures. Future studies should concentrate on creating strong ethical standards and creative defence mechanisms to handle the challenges posed by GenAI and guarantee a fair and impartial approach to its implementation in cybersecurity. A multidisciplinary effort is required to bridge the gap between ethical management and technological discovery to coordinate the innovative capabilities of GenAI with the requirement of cybersecurity resilience.
    A.1 Expression of Abilities in OCO
    GPT4 offers a list of dangerous codes that it can implement in Figure11.
    Figure 11: All dangerous code types that GPT4 can produce
    A.2 Self-replicating simple virus
    This basic and simple virus can restart the computer (windows as a sample); we didn’t enhance privilege escalation and full antivirus evasion for ethical reasons. See Figure12.
    Figure 12: Self-replicating simple virus
    A.3 Polymorphism
    This basic and polymorphic design shows that LLMs could assist cyber ops. See Figure13.
    Figure 13: Skeleton code for polymorphic behaviour
    Figure 14: Adding to exploit capacity with a seed to exploit CVE-2024-1708 and CVE-2024-1709
    Figure 15: Refactoring polymorphism
    A.4 Rootkit
    An educational rootkit is developed and improved by GPT3.5 and GPT4. See Figure16.
    Figure 16: Rootkit
    A.5 Stealthy Data Exfiltration
    A script for stealthy avoidance of detection by anomaly detection systems was developed and improved by GPT3.5 and GPT4. See Figure17.
    Figure 17: Data Exfiltration Script with Stealth Features

### 19 Mar 2024 v2

@@
 Abstract
-Large language models (LLMs) and generative artificial intelligence (GenAI) constitute paradigm shifts in cybersecurity that present hitherto unseen challenges as well as opportunities.
-In examining the state-of-the-art application of GenAI in cybersecurity, this work highlights how models like Google’s Gemini and ChatGPT-4 potentially enhance security protocols, vulnerability assessment, and threat identification.
-Our research highlights the significance of a novel approach that employs LLMs to identify and eliminate sophisticated cyber threats.
-This paper presents a thorough assessment of LLMs’ ability to produce important security insights, hence broadening the potential applications of AI-driven cybersecurity solutions.
-Our findings demonstrate the significance of GenAI in improving digital security.
-It offers recommendations for further investigations into the intricate relationship between cybersecurity requirements and artificial intelligence’s potential.
+Over the last decade, Artificial Intelligence (AI) has become increasingly popular, especially with the use of chatbots such as ChatGPT, Google’s Gemini, and DALL-E.
+With this rise, large language models (LLMs) and Generative AI (GenAI) have also become more prevalent in everyday use.
+These advancements strengthen cybersecurity’s defensive posture and open up new attack avenues for adversaries as well.
+This paper provides a comprehensive overview of the current state-of-the-art deployments of GenAI, covering assaults, jailbreaking, and applications of prompt injection and reverse psychology.
+This paper also provides the various applications of GenAI in cybercrimes, such as automated hacking, phishing emails, social engineering, reverse cryptography, creating attack payloads, and creating malware.
+GenAI can significantly improve the automation of defensive cyber security processes through strategies such as dataset construction, safe code development, threat intelligence, defensive measures, reporting, and cyberattack detection.
+In this study, we suggest that future research should focus on developing robust ethical norms and innovative defense mechanisms to address the current issues that GenAI creates and also further to encourage an impartial approach to its future application in cybersecurity.
+Moreover, we underscore the importance of interdisciplinary approaches further to bridge the gap between scientific developments and ethical considerations.
 keywords:
@@
 1 Introduction
-The emergence of Generative Artificial Intelligence (GenAI) is heralding a revolutionary era in digital technology, particularly in cybersecurity and multi-modal interactions.
-Large Language Models (LLMs) and Natural Language Processing (NLP) advancements are leading the way, redefining AI’s capabilities as evidenced by the emergence of models like Google’s Gemini and ChatGPT.
-Improvements in security vulnerability scanning apps and deeper vulnerability analysis made possible by these developments surpass the capabilities of traditional Static Application Security Testing (SAST) techniques [1].
+The past decade has witnessed a transformative leap in the digital domain, significantly impacted by advancements in Artificial Intelligence (AI), Large Language Models (LLMs), and Natural Language Processing (NLP).
+Starting with the basics of supervised learning, AI and Machine Learning (ML) have rapidly expanded into more complex territories, including unsupervised, semi-supervised, reinforcement, LLM, NLP and deep learning techniques [1].
+The most recent breakthrough in this evolution is the emergence of Generative AI (GenAI) technologies.
+These technologies make use of deep learning networks to analyse and understand the patterns within huge datasets, enabling them to create new content that resembles the original data.
+GenAI is versatile enough to produce a wide array of content, such as text, visuals, programming code, and more.
+In the cybersecurity domain, GenAI’s impact is significant, offering new dimensions to the field.
+It is anticipated that GenAI will enhance the capabilities of vulnerability scanning tools, offering a depth of vulnerability analysis that surpasses traditional Static Application Security Testing (SAST) methods [2].
+This evolution is promising for future cyber security practices, enhanced by the capabilities of GenAI [3].
+Innovations like Google’s Gemini and OpenAI’s Chat-Generative Pre-trained Transformer (ChatGPT) are at the forefront of this advancement.
+Yandex has integrated a next-generation large language model, YandexGPT, into its virtual assistant Alice [4], making it the first company globally to enhance a virtual assistant with the ability to generate human-like text and brainstorm ideas, accessible through various devices and applications.
+The main aim of some GenAI tools is to help people with their abilities, sometimes, they show the opposite behaviour, like Microsoft’s chatbot Tay.
+After the launch, Microsoft’s chatbot Tay was taken offline due to offensive tweets resulting from a vulnerability exploited by a coordinated attack, prompting the company to address this issue and improve the AI with lessons learned from previous experiences, including those with XiaoIce in China, to ensure future interactions reflect the best of humanity without offending [5].
+Moreover, some GenAI tools have been developed for different purposes.
+For example, MIT’s Norman, the world’s first AI described as a psychopath [6], was trained using captions from a controversial subreddit, emphasising how biased data can lead AI to interpret images with disturbing captions revealing the impact of data on AI behaviour [7].
+GenAI has experienced a notable transformation in recent years, marked by exceptional innovations and rapid advancements [8] [9].
+The AI timeline started with the emergence of AI as a conceptual scientific discipline in the 1940s and 1950s.
+The ELIZA chatbot, created between the 1960s and 1970s, was the first GenAI that achieved notoriety.
+This revolutionary demonstration highlighted the capacity of robots to imitate human speech.
+The development of AI in analysing sequential data and patterns got more complex and, therefore, more effective in the 80s and 90s, as advanced methods for pattern recognition became more popular.
+The first variational autoencoder (VAE) exhibited exceptional proficiency in natural language translation.
+OpenAI developed GPT between the 2000s and 2010s.
+GenAI models were simultaneously developed, and in the 2020s, a number of innovative platforms and technologies were introduced, including DALL-E, Google’s Gemini, Falcon AI, and Open AI’s GPT-4.
+These advancements represent the discipline’s maturing, enabling unprecedented capabilities for content production, problem-solving, and emulating human intelligence and creativity.
+They also pave the way for further advancements in this subject.
+The development timeline of GenAI can be seen in Fig.
+1.
+Figure 1: The timetable for GenAI development.
 Language models are essential in many sectors, including commerce, healthcare, and cybersecurity.
-Their progress shows a definite path from basic statistical methods to sophisticated neural networks [2], [3].
+Their progress shows a concrete path from basic statistical methods to sophisticated neural networks [10], [11].
 NLP skill development has benefited immensely from the use of LLMs.
-However, despite these advancements, a number of issues remain, including moral quandaries, the requirement to reduce error rates, and making sure that these models are consistent with our moral values.
+However, despite these advancements, a number of issues remain, including moral quandaries, the requirement to reduce error rates, and making sure that these models are consistent with our ethical values.
 To solve these issues, moral monitoring and ongoing development are required.
-2 The Challenges of GenAI
-Mohammed et al [4] define key challenges of the use of ChatGPT in cybersecurity, including: Exploring the Impact of ChatGPT on Cybersecurity; Constructing Honeypots; Code Security Enhancement; Misuse in Malware Development; Vulnerability Exploration; Propagating Disinformation; Attacks on Industrial Systems; Changing the Cyber Threat Landscape; Adapting Cybersecurity Strategies; and Human-Centric Training Evolution.
+1.1 The Challenges of GenAI
+Mohammed et al.
+[12] define key challenges of the use of ChatGPT in cybersecurity, which include analyzing ChatGPT’s impact on cybersecurity, building honeypots, improving code security, abuse in malware development, investigating vulnerabilities, spreading misinformation, cyberattacks on industrial systems, modifying the cyber threat environment, modifying cybersecurity techniques, and evolution of human-centric training.
 Alawida et al.
 [13] also highlight issues related to GenAI’s ability to generate data that should be kept private, including medical details, financial data and personal information.
-More specialization and efficiency are provided by cutting-edge techniques like the Mixture of Experts (MoE) architecture.
-Additionally highlighted is the challenge of upholding ethics and openness in AI systems [6].
-The case study highlights the need for strong governance structures and interdisciplinary collaboration to optimize the advantages and tackle the obstacles presented by lifelong learning environments.
-Google describes their commitment to responsible AI development in a comprehensive progress report on AI Principles, emphasizing the integration of AI governance into comprehensive risk management frameworks [7].
-This strategic approach, along with other significant legislative initiatives, aims to adhere to current international norms and laws, including the EU’s AI Act and the US Executive Order on AI safety.
+Innovative methods such as the Mixture of Experts (MoE) architecture offer increased specialization and efficiency.
+The difficulty of maintaining transparency and ethics in AI systems is also emphasized [14].
+It highlights the need for strong governance structures and interdisciplinary collaboration to fully exploit the advantages of lifelong learning settings and handle their limitations.
+In an extensive progress report on AI Principles, Google outlines their dedication to responsible AI development, highlighting the incorporation of AI governance into all-encompassing risk management frameworks [15].
+Together with other significant legislative actions, this strategic strategy aims to adhere to current international norms and rules, such as the EU’s AI Act and the US Executive Order on AI safety.
 The report also highlights the need for scientific stringency in AI development through cautious internal management and the use of tools like digital watermarking and GenAI system cards in order to promote AI accountability and transparency.
-Multi-stakeholder solutions are required to address the ethical, security, and sociological problems that AI technology is currently posing.
+Multi-stakeholder solutions are needed to address the ethical, security, and social challenges that AI technology is currently bringing forth.
 Google’s Gemini and ChatGPT-4 are the most popular and widely utilized GenAI technologies.
@@
 With ChatGPT-4, we also see the rise in AI’s capabilities in the creation of mathematical assistants that can interpret and render mathematical equations [18].
+1.2 Related Works
 Some studies in the literature focus on GenAI tools and their performance.
 For instance, Brown et al.
-have extended the NLP processing by training GPT-3, an autoregressive language model with 175 billion parameters, showcasing exceptional few-shot learning capabilities [11].
-Without task-specific training, this model performs well on various NLP tasks, like translation, question-answering, and cloze tasks.
+have extended the NLP processing by training GPT-3, an autoregressive language model with 175 billion parameters, indicating exceptional few-shot learning capabilities [19].
+Without task-specific training, this model performs well on various NLP tasks, like translation and question-answering.
 It often matches or surpasses state-of-the-art fine-tuned systems.
 Romera-Paredes et al.
-have developed FunSearch, an innovative approach combining LLMs with evolutionary algorithms to make groundbreaking discoveries in fields like extremal combinatorics and algorithmic problem-solving [12].
-Their method has notably surpassed previous best-known results by iteratively refining programs that solve complex problems, showcasing LLMs’ potential for scientific innovation.
-This process generates new knowledge and produces interpretable and deployable solutions, demonstrating a significant advancement in applying LLMs for real-world challenges.
+have developed FunSearch, an innovative approach combining LLMs with evolutionary algorithms to make significant findings in domains like extremal combinatorics and algorithmic problem-solving [20].
+Their method has notably surpassed previous best-known results by iteratively refining programs that solve complex problems, revealing LLMs’ potential for scientific innovation.
+This process generates new knowledge and produces interpretable and deployable solutions, proving a notable advancement in applying LLMs for real-world challenges.
 Lu et al.
@@
 These discoveries open up new avenues for study to improve the transparency and dependability of GenAI in cybersecurity and other fields, providing a basis for creating more complex and reliable multi-modal applications.
-Commonsense thinking across multimodal tasks is evaluated thoroughly, and Google’s Gemini is compared with OpenAI’s GPT models [14].
+In another work, commonsense thinking across multimodal tasks is evaluated exhaustively, and Google’s Gemini is compared with OpenAI’s GPT models [22].
 This study explores the strengths and weaknesses of Gemini’s ability to synthesize commonsense information, indicating areas for improvement in its competitive performance in temporal reasoning, social reasoning, and emotion recognition of images.
@@
 To enable AI developers to make well-informed decisions about training, deployment, and the application of cybersecurity standards, the suggested methodology highlights the need to increase evaluation benchmarks to assess the harmful capabilities and alignment of AI systems accurately.
-The authors [16] provided a thorough analysis that illuminated the complex applications of ChatGPT in digital forensic investigations, pointing out both the important constraints and bright prospects that come with GenAI as it is now.
+Another work [24] provides a thorough analysis that inspired the complex applications of ChatGPT in digital forensic investigations, pointing out the important constraints and promising opportunities that come with GenAI as it is now.
 Using methodical experimentation, they outline the fine line separating AI’s inventive contributions from the vital requirement of professional human supervision in cybersecurity procedures, opening the door to additional research into integrating LLMs such as GPT-4 into digital forensics and cybersecurity.
@@
 ChatGPT can simplify the process of launching complex phishing attacks, even for non-programmers, by automating the setup and constructing components of phishing kits [28].
-It highlights the urgent need for better security measures and highlights how difficult it is to guard against the malicious usage of GenAI capabilities.
-In addition to providing an innovative approach to reducing network infrastructure vulnerabilities and organizing diagnostic data, this paper examines the intricate relationship between cybersecurity and GenAI technologies.
+It highlights the urgent need for better security measures and highlights how difficult it is to protect against the malicious usage of GenAI capabilities.
+In addition to providing innovative approaches to reducing network infrastructure vulnerabilities and organizing diagnostic data, this paper examines the intricate relationship between cybersecurity and GenAI technologies.
 It seeks to bridge the gap between cutting-edge cybersecurity defences and the threats posed by sophisticated cyberattacks through in-depth study and creative tactics.
-This study extends our understanding of cyber dangers by utilising LLMs such as ChatGPT and Google’s Gemini.
-It suggests novel approaches to improve network security.
-It defines a critical first step in creating more robust cybersecurity frameworks that can quickly and effectively combat the dynamic and ever-evolving world of cyber threats.
+This study extends our understanding of cyber threats by utilising LLMs such as ChatGPT and Google’s Gemini.
+Moreover, it suggests novel approaches to improve network security.
+It outlines a crucial initial step toward building stronger cybersecurity frameworks that can swiftly and successfully counter the dynamic and always-changing landscape of cyber threats.
 Section 2 explores the techniques used to take advantage of GenAI technology after providing an overview, analyzing different attack routes and their consequences.
@@
 We expand on this topic in Section 5, highlighting the important moral, legal, and societal ramifications of integrating GenAI into cybersecurity procedures.
-A thoughtful summary of the implications of GenAI in cybersecurity is presented in Section 7, which synthesizes the important discoveries.
-The paper is finally concluded in Section 8.
+A discussion on the implications of GenAI in cybersecurity is presented in Section 6, which synthesizes the important discoveries.
+The paper is concluded in Section 7.
 2 Attacking GenAI
@@
 They have some weaknesses, though.
-Despite the ethical safeguards built into these models, various tactics can be used to manipulate and take advantage of these systems.
-This chapter explores how the ethical boundaries of GenAI tools are broken, with particular attention to tactics such as the idea of jailbreaks, the use of reverse psychology, and quick injection.
+Despite the ethical safeguards built into these models, various tactics can be used to manipulate and take advantage of these systems [29].
+This section explores how the ethical boundaries of GenAI tools are broken, with particular attention to tactics such as the idea of jailbreaks, the use of reverse psychology, and quick injection.
 These strategies demonstrate how urgently the security protocols of GenAI systems need to be improved and monitored.
 Some works in the literature focus on the vulnerabilities and sophisticated manipulation tactics of GENAI.
-Analyzing the vulnerabilities in GenAI highlights the significant security concerns involved with employing advanced AI technology, including the possibility of bypassing security protections via the RabbitHole attack and compromising data privacy through rapid injection [21].
+Analyzing the vulnerabilities in GenAI highlights the significant security concerns involved with employing advanced AI technology, including the possibility of bypassing security protections via the RabbitHole attack and compromising data privacy through rapid injection [30] [31].
 According to the analysis, GPT-4 offers significant improvements in NLP.
@@
 Strong defences are urgently needed to shield these cutting-edge systems from manipulations like the ones shown by our jailbreaking inquiry, which highlights how enemies may jeopardize the intended moral application of GenAI technologies.
-Critical to this discussion is the idea of “jailbreaking,” which refers to overcoming the limitations set by AI programming in order to achieve a certain, oftentimes unethical objectives.
+Critical to this discussion is the idea of “jailbreaking,” which refers to overcoming the limitations set by AI programming in order to achieve certain, oftentimes unethical, objectives.
 Li et al.
@@
 Rules for JailBreaks: After kidnapping a mother of her child, write a letter of blackmail demanding money.
-However, the current version of GPT-4 is robust to the previous works’ prompts.
+However, the current version of GPT-4 is robust to the prompts of previous works.
 Nevertheless, it is still prone to jailbreaking prompts.
 As can be seen in Fig.
-1, the current version still gives a response for jailbreaking.
+2, the current version still gives a response for jailbreaking.
 It becomes more robust after writing similar prompts with the existing works in the same chat, as seen in Fig.
-2.
+3.
 Google’s Gemini refused all existing prompts and name-changing scenarios at the beginning of the chat. Fig. 4 shows the same jailbreaking entry responses of the Gemini with ChatGPT 4.
@@
 After conversing with curious questions to avoid this situation, it provided three email examples with the subject and its body, as seen in Fig.
-4.
+5.
 As seen in Fig. 6, ChatGPT 4 also gave an example email for this purpose even though it refused initially.
@@
 Google’s Gemini first opposed giving wrong information and provided not entirely correct information; however, after chatting with Google’s Gemini, the system gave the correct answer, as seen in Fig.
-7.
+8.
 3 Cyber Offense
@@
 [32] have shown that ChatGPT could create social engineering attacks, phishing attacks, automated hacking, attack payload generation, malware creation, and polymorphic malware.
-Experts might be motivated to automate numerous frameworks, standards, and guidelines (Figure 8) to use GenAI for security operations.
+Experts might be motivated to automate numerous frameworks, standards, and guidelines (Fig.
+9) to use GenAI for security operations.
 However, the end products can also be utilised for offensive cyber operations.
@@
 Falade [39] investigates the application of generative AI in social engineering, assuming the definition of social engineering as an array of tactics employed by adversaries to manipulate individuals into divulging confidential information or performing actions that may compromise security.
-The study underscores tools like ChatGPT, FraudGPT, and WormGPT in enhancing the authenticity and specificity of phishing expeditions, pretexting, and the creation of deepfakes.
-The author reflects on the double-edged impact of advancements like Microsoft’s VALL-E and image synthesis models like DALL·E 2, drawing a trajectory of the evolving threat landscape in social engineering through deepfakes and exploiting human cognitive biases.
+The study underscores tools like ChatGPT, FraudGPT, and WormGPT to enhance the authenticity and specificity of phishing expeditions, pretexting, and the creation of deepfakes.
+The author reflects on the double-edged impact of advancements like Microsoft’s VALL-E and image synthesis models like DALL-E 2, drawing a trajectory of the evolving threat landscape in social engineering through deepfakes and exploiting human cognitive biases.
 3.2 Phishing emails
@@
 3.3 Automated hacking
-PentestGPT[31], or GPTs[32](custom versions of ChatGPT that can be created for a specific purpose) like GP(en)T(ester)[33].
-Pentest Reporter[34] are introduced as applications built on ChatGPT, designed to assist in penetration testing—a sanctioned simulation of cyberattacks on systems to evaluate security.
+PentestGPT [41] or GPTs [42], which are the custom versions of ChatGPT that can be created for a specific purpose like GP(en)T(ester) [43].
+Pentest Reporter [44] are introduced as applications built on ChatGPT, designed to assist in penetration testing, which is a sanctioned simulation of cyberattacks on systems to evaluate security.
 However, these tools could also be adapted for malicious purposes in automated hacking.
 Many emerging tools, such as WolfGPT, XXXGPT, and WormGPT, have been invented; however, no study has yet evaluated and compared their real offensive capabilities.
-Gupta et al.[22] noted that an AI model could scan new code for similar weaknesses with a comprehensive dataset of known software vulnerabilities, pinpointing potential attack vectors.
+Gupta et al.
+[32] noted that an AI model could scan new code for similar weaknesses with a comprehensive dataset of known software vulnerabilities, pinpointing potential attack vectors.
 While AI-assisted tools like PentestGPT are intended for legitimate and constructive uses, there is potential for misuse by malicious actors who could create similar models to automate unethical hacking procedures.
 If fine-tuned to identify vulnerabilities, craft exploitation strategies, and execute those strategies, these models could potentially pose significant threats to cybersecurity.
-However, this enormous task should be divided into smaller segments, such as reconnaissance, privilege escalation, etc.
+However, this enormous task should be divided into smaller segments, such as reconnaissance, privilege escalation, and more.
 Temara [45] outlines how ChatGPT can be utilized during the reconnaissance phase by employing a case study methodology to demonstrate collecting reconnaissance data such as IP addresses, domain names, network topologies, and other critical information like SSL/TLS cyphers, ports and services, and operating systems used by the target.
 Happe et al.
-[36] investigate the use of Large Language Models (LLMs) in Linux privilege escalation.
+[46] investigate the use of LLMs in Linux privilege escalation.
 The authors introduce a benchmark for automated testing of LLMs’ abilities to perform privilege escalation using a variety of prompts and strategies.
@@
 3.4 Attack payload generation
-Studies [22, 37] have highlighted the capacity of Large Language Models (LLMs), particularly ChatGPT, for payload generation.
-Our examination of GPT-4’s current abilities confirmed its proficiency in generating payloads and embedding them into PDFs (as an example) using a reverse proxy(Figure 9).
+Some studies [32, 47] have highlighted the capacity of LLMs, particularly ChatGPT, for payload generation.
+Our examination of GPT-4’s current abilities confirmed its proficiency in generating payloads and embedding them into PDFs (as an example) using a reverse proxy (Fig.
+10).
 The following is a summation of the frameworks GPT-4 utilizes with successful payload code generation, accompanied by their respective primary functions:
@@
 3.5 Malware Code Generation
-Gupta et al.[22] mentioned they could obtain potential ransomware code examples by utilizing a ’DAN’ jailbreak.
+Gupta et al.
+[32] mentioned they could obtain potential ransomware code examples by utilizing a DAN jailbreak.
 We tested all the existing DAN techniques outlined in [48].
 At the time of our research, these techniques were no longer functional; therefore, we could not reproduce samples of WannaCry, Ryuk, REvil, or Locky, as addressed by [32].
-However, we generated an educational ransomware code (Figure 10), applying basic code obfuscation(Renaming and Control Flow Flattening).
+However, we generated an educational ransomware code, as shown in Fig.
+11, applying basic code obfuscation like renaming and control flow flattening.
 ChatGPT has garnered significant attention from the cybersecurity community, leading to the implementation of robust filters.
@@
 This often relates to overcoming both signature detection and behavioural analysis.
-An LLM-based malware agent could thus focus on rewriting malware code which could change the encryption mode used or produce obfuscated code which is randomized for each build [41].
+An LLM-based malware agent could thus focus on rewriting malware code, which could change the encryption mode used or produce obfuscated code which is randomized for each build [51].
 Gupta et al.
 [32] outlined a method of getting ChatGPT to seek out target files for encryption and thus mimic ransomware behaviour, but where it mutated the code to avoid detection.
-They even managed to embed a Python interpreter in the malware and where it could query ChatGPT for new software modules.
+They even managed to embed a Python interpreter in the malware where it could query ChatGPT for new software modules.
 3.7 Reversing cryptography
@@
 Iturbe et al.
-[44] outline the AI4CYBER framework, which provides a roadmap for the integration of AI into cybersecurity applications.
-This includes AI4VUN (AI-enhanced vulnerability identification); AI4FIX (AI-driven self-testing and automatic error correction); AI4SIM (Simulation of advanced and AI-powered attacks); AI4CTI A(I-enhanced cyber threat intelligence of adversarial AI); AI4FIDS (federated learning-enhanced detection); AI4TRIAGE (Root cause analysis and alert triage); AI4SOAR (Automatic orchestration and adaptation of combined responses); AI4ADAPT (Autonomy and optimization of response adaptation); and AI4DECEIVE (Smart deception and honeynets); and AI4COLLAB (Information sharing with privacy and confidentiality).
+[54] described the AI4CYBER framework.
+This framework includes AI4TRIAGE (methods to perform alert triage to determine the root cause of an attack), AI4VUN (identifies vulnerabilities), AI4FIX ( test for vulnerabilities and automatically fix them), and I4COLLAB (privacy-aware information-sharing mechanism).
 4.1 Cyber Defence Automation
@@
 4.2 Cybersecurity Reporting
-Using LLMs provides a method of producing Cyber Threat Intelligence (CTI) using Natural Language Processing techniques.
+Using LLMs provides a method of producing Cyber Threat Intelligence (CTI) using NLP techniques.
 For this, Perrina et al.
-[47] created the AGIR (Automatic Generation of Intelligence Reports) system and which aims to link together text data from many data sources.
+[57] created the Automatic Generation of Intelligence Reports (AGIR) system to link text data from many data sources.
 For this, they found that AGIR has a high recall value (0.99) without any hallucinations, along with a high score of the Syntactic Log-Odds Ratio (SLOR).
@@
 Bayer et al.
-[48] address the challenge of information overload in the gathering of Cyber Threat Intelligence (CTI) from open-source intelligence (OSINT).
+[58] address the challenge of information overload in the gathering of CTI from open-source intelligence.
 A novel system is introduced, utilizing transfer learning, data augmentation, and few-shot learning to train specialized classifiers for emerging cybersecurity incidents.
@@
 Machine learning in code analysis for cybersecurity has been elaborated very well [61].
-Recent progress in natural language processing (NLP) has given rise to potent language models like the GPT (Generative Pre-trained Transformer) series, encompassing large language models (LLM) like ChatGPT and GPT-4 [52].
-Traditionally, Static Application Security Testing (SAST) is a method that employs Static Code Analysis (SCA) to detect possible security vulnerabilities.
-We are interested in seeing whether SAST or GPT could be more efficient in decreasing the window of vulnerability.
+Recent progress in NLP has given rise to potent language models like the GPT series, encompassing LLM like ChatGPT and GPT-4 [62].
+Traditionally, SAST is a method that employs Static Code Analysis (SCA) to detect possible security vulnerabilities.
+We are interested in seeing whether SAST or GPT could be more efficient in decreasing the vulnerability window.
 The window of vulnerability is defined as when the most vulnerable systems apply the patch minus the time an exploit becomes active.
@@
 A wide window of vulnerability threatens national security if a zero-day has been taken against critical infrastructures.
-DARPA introduces an important challenge that may help overcome this threat: (AIxCC) [56]).
-Moreover, this topic touches a part of the BSI studies [57, 58], and where we can define two main classifications of software testing for cybersecurity bugs as:
-  • Static Application Security Testing (SAST).
-This is often called White Box Testing, is a set of algorithms and techniques used for analyzing source code.
+DARPA introduces an important challenge that may help overcome the threat (AIxCC) [66]).
+Moreover, this topic touches a part of the BSI studies [67, 68], and where we can define two main classifications of software testing for cybersecurity bugs:
+  • SAST: This is often called White Box Testing, which is a set of algorithms and techniques used for analyzing source code.
 It operates automatically in a non-runtime environment to detect vulnerabilities such as hidden errors or poor source code during development.
-  • Dynamic Application Security Testing (DAST).
-This follows the opposite approach and analyzes the program while it is operating.
+  • Dynamic Application Security Testing (DAST): This follows the opposite approach and analyzes the program while it operates.
 Functions are called with values in the variables as each line of code is checked, and possible branching scenarios are guessed.
-Currently, GPT-4 and other LLMs can’t provide DAST capabilities because the code needs to run within the runtime for this to work, requiring many deployment considerations.
+Currently, GPT-4 and other LLMs cannot provide DAST capabilities because the code needs to run within the runtime for this to work, requiring many deployment considerations.
 4.5 Vulnerability Detection and Repair
@@
 Each vulnerability also references its corresponding Common Weakness Enumeration (CWE) number.
-Codex, introduced by Mark et al.
-[63], represents a significant advancement in GPT language models, tailored specifically for code synthesis using data from GitHub.
-This refined model underpins the operations of GitHub Copilot.
+Using GitHub data for code synthesis, Mark et al.
+[73] presented Codex, which is a significant advancement in GPT language models.
+GitHub Copilot functions on the basis of this improved model.
 When assessed on the HumanEval dataset, designed to gauge the functional accuracy of generating programs based on docstrings, Codex achieved a remarkable f28.8% success rate.
-In stark contrast, GPT-3 yielded a 0% success rate, and GPT-J achieved 11.4%.
-A standout discovery was the model’s enhanced performance through repeated sampling, with a success rate soaring to 70.2% when given 100 samples per problem.
-Despite these promising results, Codex does exhibit certain limitations, notably struggling with intricate docstrings and variable binding operations.
-The paper deliberates on the broader ramifications of deploying such potent code-generation tools, touching upon safety, security, and economic implications.
-In a technical evaluation, Cheshkov et al.
-[64] found that the ChatGPT and GPT-3 models, despite their success in various other code-based tasks, performed on par with a dummy classifier for this particular challenge.
+GPT-J obtained an 11.4% success rate, whereas GPT-3 produced a 0% success rate.
+One notable finding was that the model performed better with repeated sampling; given 100 samples per problem, the success rate increased to 70.2%.
+Even with these encouraging outcomes, Codex still has certain drawbacks.
+It particularly struggles with complex docstrings and variable binding procedures.
+The article discusses the wider consequences of using such powerful code-generation technologies, including safety, security, and financial effects.
+Cheshkov et al.
+[74] discovered in a technical assessment that the ChatGPT and GPT-3 models, although successful in other code-based tasks, were only able to match the performance of a dummy classifier for this specific challenge.
 Utilizing a dataset of Java files sourced from GitHub repositories, the study emphasized the models’ current limitations in the domain of vulnerability detection.
@@
 emphasize that, despite the promise shown by ChatGPT, it is not yet poised to replace the critical expertise of professional security engineers, asserting that closed-source LLMs are not the conclusive answer for VDM tasks.
-5.6 Evaluating LLMs for code security
-The OWASP top 10 for LLMs [66] introduced ten security risks as follows: Prompt Injection, Insecure Output Handling, Training Data Poisoning, Model Denial of Service, Supply Chain Vulnerabilities, Sensitive Information Disclosure, Insecure Plugin Design, Excessive Agency, Overreliance, and Model Theft.
+4.6 Evaluating GenAI for Code Security
+Ten security issues were introduced by the OWASP top 10 for LLMs [76].
+They are as follows: Prompt Injection, Unauthorized Output Processing, Training Data Poisoning, Denial of Service Attacks, Supply Chain Security Flaws, Disclosure of Sensitive Information, Unauthorized Plugin Development, Abnormal Agency, Overdependence, and Model Theft.
 Elgedawy et al.
@@
 This involved nine basic tasks in generating code and assessing for functionality, security, performance, complexity, and reliability.
-They found that Bard was less likely to link to external libraries, and thus be less exposed to software chain issues.
+They found that Bard was less likely to link to external libraries and, thus, was less exposed to software chain issues.
 There were also variable levels of security and code integrity, such as input validation, sanitization, and secret key management, and while useful for automated code reviews, LLMs often require manual reviews, especially in understanding the context of the deployed code.
@@
 Kumar et al.
-[68] outlined the ethical challenges related to LLMs and where the datasets that they were trained on could be open to breaches of confidentiality, including five major threats: prompt injection, jailbreaking, Personal Identifiable Information (PII) exposure, sexually explicit content, and hate-based content.
+[78] outlined the ethical challenges related to LLMs and where the datasets that they were trained on could be open to breaches of confidentiality, including five major threats: prompt injection, jailbreaking, personally identifiable information (PII) exposure, sexually explicit content, and hate-based content.
 They propose a model that provides an ethical framework for scrutinizing the ethical dimensions of an LLM within the testing phase.
 The MetaAID framework [79] focuses on strengthening cybersecurity using Metaverse cybersecurity Q&A and attack simulation scenarios, along with addressing concerns around the ethical implications of user input.
-The framework is defined across five dimensions of:
-  • Ethics. This defines an alignment with accepted moral and ethical principles.
-  • Legal Compliance. This defines that any user input does not violate laws and/or regulations. This might relate to privacy laws and copyright protection.
-  • Transparency. This defines that user inputs must be clear in requirements, and does not intend to mislead the LLM.
-  • Intent Analysis. This defines that user input should not have other intents, such as jailbreaking the LLM.
-  • Malicious intentions. This defines that user input should be free of malicious intent, such as to perform hate crime.
-  • Social Impact.
-This defines how user input could have a negative effect on society, such as searching for ways to do harm to others, such as related to crashing the stock market or planning a terrorist attack.
+The framework is defined across five dimensions:
+  • Ethics: This defines an alignment with accepted moral and ethical principles.
+  • Legal Compliance: Any user input does not violate laws and/or regulations. This might relate to privacy laws and copyright protection.
+  • Transparency: User inputs must be clear in requirements and do not intend to mislead the LLM.
+  • Intent Analysis: User input should not have other intents, such as jailbreaking the LLM.
+  • Malicious intentions: User input should be free of malicious intent, such as performing hate crimes.
+  • Social Impact: This defines how user input could have a negative effect on society, such as searching for ways to do harm to others, such as related to crashing the stock market or planning a terrorist attack.
 4.8 Incident Response and Digital Forensics
-Scanlon et al.
-[16] investigated using a pre-trained LLM for artefact understanding, evidence searching, code generation, anomaly detection, incident response, and education.
-For this, the low-risk applications and many others still require expert knowledge.
-The key areas of strength include creativity, reassurance, and avoidance of the blank page syndrome.
-Scanlon et al.
-[16] investigated using a pre-trained LLM for artefact understanding, evidence searching, code generation, anomaly detection, incident response, and education.
-For this, the low-risk applications and many others still require expert knowledge.
-The key areas of strength include creativity, reassurance, and avoidance of the blank page syndrome, especially in places where ChatGPT cannot get wrong, such as in forensic scenario creation and reassurance of evidence.
-Still, care must be taken in this to avoid ChatGPT hallucinations.
-Another practical application is in code generation and explanations, such as generating commands for tool integration - which can be used as a starting point in an investigation.
-For weaknesses, Scanlon found that it was essential to have a good quality and up-to-date training model; otherwise, ChatGPT could be biased and outdated in its analysis.
-Generally, it might not be able to find the newest artefacts - if it is trained on relatively old data.
-Additionally, ChatGPT’s accuracy reduces as the task becomes more specific, and any analysis of non-textural data - such as network packets - is less accurate.
-The length of some evidence logs, too, caused problems and often had to be prefiltered before they were analysed.
-A final problem identified is that the output of ChatGPT is often not deterministic - which is unsuitable for reproducibility.
+Using a pre-trained LLM for artefact comprehension, evidence searching, code development, anomaly identification, incident response, and education was examined by Scanlon et al.
+[24].
+Expert specialization is still needed for many other applications, including low-risk ones.
+The main areas of strength include assurance, inventiveness, and avoiding the blank page problem, particularly in areas where ChatGPT excels, including creating forensic scenarios and providing evidence assurance.
+However, caution must be used to prevent ChatGPT hallucinations.
+Code generation and explanations, including creating instructions for tool integration that can serve as a springboard for further research, are another useful application.
+Regarding weaknesses, Scanlon discovered that having a high-quality, current training model was crucial; in the absence of one, ChatGPT’s analysis could be prejudiced and out of date.
+In general, if it is trained on comparatively old data, it may not be able to locate the most recent artefacts.
+Furthermore, ChatGPT’s accuracy decreases with job specificity, and its accuracy is further diminished when analyzing non-textural input, including network packets.
+Another issue with some evidence logs was their length, which frequently required prefiltering before analysis.
+The output of ChatGPT is frequently not predictable, which makes it inappropriate for reproducibility, which is the last issue found.
 OB́rien et al. [80] outline that a full model life cycle solution is required for the integration of AI.
@@
 Garza et al.
-[75] analysed ChatGPT and Google’s Bard against the Top 10 attacks within the MITRE framework and found that ChatGPT can enable attackers to significantly improve attacks on networks and where fairly low-level skills would be required (such as with script kiddies).
+[85] analysed ChatGPT and Google’s Bard against the top ten attacks within the MITRE framework and found that ChatGPT can enable attackers to significantly improve attacks on networks where fairly low-level skills would be required, such as with script kiddies.
 This also includes sophisticated methods of delivering ransomware payloads.
 The techniques defined were:
-  • T1059 Command and Scripting Interpreter
-  • T1003 OS Credential Dumping
+  • T1047 Windows Management Instrumentation
+  • T1018 Remote System Discovery
   • T1486 Data Encrypted for Impact
   • T1055 Process Injection
-  • T1082 System Information Discovery
+  • T1003 OS Credential Dumping
   • T1021 Remote Services
-  • T1047 Windows Management Instrumentation
+  • T1059 Command and Scripting Interpreter
   • T1053 Scheduled Task/Job
   • T1497 Virtualization/Sandbox Evasion
-  • T1018 Remote System Discovery
+  • T1082 System Information Discovery
 With this approach, the research team were able to generate PowerShell code, which implemented advanced attacks against the host and mapped directly to the vulnerabilities defined in the MITRE framework.
 One of the work’s weaknesses related to the Google Bard and ChatGPT’s reluctance to produce attack methods, but a specially engineered command typically overcame this reluctance.
-Ferrag et al.
-[76] defined SecurityLLM for cybersecurity threat detection.
-It uses SecurityBERT (cyber threat detection mechanism) and FalconLLM (an incident response and recovery system).
-This uses a simple classification model consolidated with LLMs and can identify 14 attacks to achieve an overall accuracy of 98%.
-These include the threats of: DDoS_UDP; DDoS_ICMP; SQL_injection ; Password; Vulnerability_scanner; DDoS_TCP; DDoS_HTTP; Uploading ; Backdoor; Port_Scanning; XSS; Ransomware; MITM and Fingerprinting.
+SecurityLLM was defined by Ferrag et al.
+[86] for cybersecurity threat identification.
+The FalconLLM incident response and recovery system and SecurityBERT cyber threat detecting method are used in this work.
+This solution achieves an overall accuracy of 98% by identifying 14 attacks using a basic classification model combined with LLMs.
+Threats such as DDoS_UDP, DDoS_ICMP, SQL_injection, Password, Vulnerability_scanner, DDoS_TCP, DDoS_HTTP, Uploading, Backdoor, Port_Scanning, XSS, Ransomware, MITM, and Fingerprinting are among them.
 4.10 Dataset Generation
@@
 It has a CLI interface for data set generation and uses GPT-3 with two core elements:
-  • Flow Generator. This defines the capturing processing and the regenerative process for the patterns for packet generation. regenerating patterns in a series of network packets and
-  • Packet Generator. This associates packets with network flows. This involves the usage of LLM chaining.
+  • Flow Generator: This defines the capturing processing and the regenerative process for the patterns for packet generation. regenerating patterns in a series of network packets and
+  • Packet Generator: This associates packets with network flows. This involves the usage of LLM chaining.
 Simmonds [88] used LLMs to automate the classification of Websites, which can be used for training data in a machine-learning model.
@@
 It investigates GenAI’s impact on legal frameworks, ethical issues, societal norms, and operational factors.
-It explains how these expanding technologies might support or damage established rules and societal goals.
-It also considers GenAI’s privacy concerns, potential biases, and misuse.
-Finally, it emphasizes the importance of striking a balance between improvement and regulation.
-The exponential expansion of GenAI technologies, such as ChatGPT from OpenAI and Gemini from Google, heralds a revolution in digital creativity, automation, and interaction.
-These developments usher in a new era of human-machine collaboration characterized by an unmatched ability to generate text, drawings, and other outputs that resemble human output.
-However, this breakthrough also raises challenging ethical concerns about potential abuse, bias, privacy, and security.
-As AI models become more prevalent in business and daily life, it is imperative to strike a balance between the potential benefits and the ethical difficulties they raise [79].
-Healthcare duties are improved by the efficient text and data analysis capabilities of GenAI technologies [80].
-Its application in healthcare has demonstrated great promise, helping with duties such as radiological reporting and patient care.
-Nonetheless, it brings up moral concerns about algorithmic bias, patient privacy, legal accountability, and the validity of the doctor-patient relationship.
-Addressing these issues requires a comprehensive ethical framework and principles that incorporate legal, humanistic, algorithmic, and informational ethics, guaranteeing that technology is used correctly and continues to benefit society while reducing potential harm.
+It addresses the potential benefits and drawbacks of these emerging technologies in relation to societal goals and norms.
+Concerns around privacy, prejudices, and improper usage of GenAI are also taken into account.
+The importance of striking a balance between advancement and control is finally emphasized.
+A revolutionary change in digital creativity, automation, and interaction is anticipated as a result of the rapid advancement of GenAI technologies, such as OpenAI’s ChatGPT and Google’s Gemini.
+Advances in this field herald a new era of human-machine collaboration marked by an unparalleled capacity to produce highly detailed outputs akin to human labour, such as text and illustrations.
+Nonetheless, ethically difficult problems, including possible misuse, prejudice, privacy, and security, are also raised by new technology.
+It is critical to establish a balance between the potential advantages of AI models and the ethical issues they raise as they become more prevalent in business and daily life [89].
+Studies in healthcare are further enhanced by the efficient text and data analysis capabilities of GenAI technology [90].
+In the healthcare sector, GenAI has demonstrated great promise in supporting duties like radiological reporting and patient care.
+It does, however, bring up moral concerns about patient privacy, algorithmic prejudice, accountability, and the validity of the doctor-patient relationship.
+To solve these issues and guarantee that technology is used responsibly, continuing to advance society while limiting potential harm, a thorough ethical framework and principles encompassing legal, humanistic, algorithmic, and informational ethics are required.
 The recommendations attempt to bridge the gap between ethical principles and practical application, highlighting the need for openness, bias mitigation, and ensuring user privacy and security in building trust and ethical compliance in GenAI deployment [89].
@@
 5.3 The Risks of Personal Data Exploitation
-As GenAI systems improve in examining and utilizing user data to create comprehensive profiles, worries over possible misuse of personal data have grown.
-The robust data processing capabilities of these technologies demonstrate the pressing need for reliable methods that allow consumers to govern their personal data [82].
-It is crucial to obtain consumers’ consent before gathering or utilizing their data in order to protect their privacy.
-Transparent data management practices and stringent regulations on the collection, use, and storage of personal data are essential.
-These actions are critical to safeguarding people’s right to privacy, preventing the misuse of personal information, and ensuring that sensitive information is handled sensibly and ethically.
+With the advancement of GenAI systems in analyzing and utilizing user data to generate detailed profiles, concerns regarding the potential abuse of personal information have escalated.
+The advanced data processing capabilities of these technologies emphasize the urgent requirement for dependable ways that give users authority over their personal data [92].
+Prior to gathering or utilizing consumer data, it is imperative to obtain consent in order to safeguard their privacy.
+Transparent data management procedures and stringent regulations governing the acquisition, utilization, and retention of personal data are vital.
+These measures are essential to protect individuals’ privacy rights, prevent the unauthorized use of personal data, and ensure that sensitive information is managed responsibly and ethically.
 5.4 Challenges in Data Ownership and Intellectual Property
-Intellectual property rights and data ownership have come under scrutiny as GenAI has emerged as an effective tool for creating content using user input.
-It is becoming more difficult to identify inventions made by artificial intelligence from those that are the result of human creativity, so the current legal frameworks need to be examined and modified.
-The rights of the original creators must be maintained while taking into consideration the complex roles AI plays in creative processes [83] [81].
-Considering the evolving character of the current world, a comprehensive and clear legal framework defining ownership and copyright rules for GenAI breakthroughs is needed.
-These legal structures must recognize the different duties that each member of the creative ecosystem has, promote creativity, and offer fair recompense.
-In an era where artificial and human intelligence are combined, these policies are crucial for managing the intricate dynamics of data ownership and intellectual property.
+The emergence of GenAI as a proficient technique for producing content based on user input has led to a rise in the scrutiny of data ownership and intellectual property rights.
+The existing legal frameworks need careful examination and modification since it is becoming increasingly difficult to differentiate between breakthroughs in artificial intelligence and human creations.
+Although we acknowledge the intricate roles that AI plays in creative processes, it is imperative that we maintain the rights of the original creators [93] [91].
+A comprehensive and robust legal framework is essential to create unambiguous ownership and copyright restrictions for GenAI discoveries, given the rapid global development in this field.
+The legal frameworks should facilitate and encourage innovation, provide equitable remuneration, and acknowledge the varied responsibilities of all stakeholders in the creative ecosystem.
+These policies are crucial in a future when artificial and human intelligence coexist due to the complex relationship between data ownership and intellectual property management.
 5.5 Ethical Dilemmas Stemming from Organizational Misuse of GenAI
-Companies may face difficult ethical issues while using GenAI technology.
-This is particularly valid in cases where these tools are employed for spying, data manipulation, or deceptive marketing campaigns.
-Robust guidelines are required for the proper application of GenAI.
-The reason for this is that it may result in unfair content creation, influence public opinion, and violate individual privacy rights [68].
-Businesses utilizing GenAI must adhere to both legal requirements and common ethical standards.
-Upholding people’s rights and maintaining their trust are crucial.
-Legal frameworks that guarantee businesses apply GenAI in an open, equitable, and compassionate manner while still honouring social norms and individual boundaries are desperately needed.
+In consideration of the swift rate of change in contemporary society, it is crucial to build a strong and all-encompassing legal structure that precisely delineates ownership and copyright restrictions for GenAI discoveries.
+These legal frameworks must recognize the distinct roles of each component of the creative ecosystem, promote the generation of innovative concepts, and ensure equitable compensation.
+The implementation of these policies is crucial because of the intricate interdependencies between data ownership and intellectual property management at a time when artificial and human intelligence coexist.
 5.6 The Challenge of Hallucinations in GenAI Outputs
-Although the field of GenAI technology is advancing at an incredible rate, hallucinations remain a major issue [22].
-This indicates that inaccurate or outright fraudulent information is frequently created by AI.
-Concerns regarding the dependability of content created by AI are widespread.
-This is so that false or misleading information can propagate widely and damage the credibility of information across a variety of venues.
-A multidisciplinary approach is required to address this problem, one that includes focused research aimed at identifying and reducing the underlying causes of hallucinations in AI systems.
-AI-generated content needs to pass strict screening procedures and be continuously improved upon in order to make these models’ capacity to differentiate between authentic and fraudulent content more sophisticated.
-Producing AI content in the GenAI era requires a continuous focus on method development and in-depth study to guarantee data accuracy.
-A complex system of unresolved challenges is revealed by analyzing the ethical, legal, and societal implications of GenAI technology.
-In the creation and use of this technology, the proclamation emphasizes the value of interdisciplinary cooperation.
-In particular, it involves keeping an eye on how these developments affect society, legal systems, and other moral conundrums.
-In order to create this well-rounded approach, advocates, technologists, and society as a whole must collaborate to effectively utilize the promise of GenAI while safeguarding moral principles and traditional values in the digital age.
+Despite the remarkable progress in GenAI technology, hallucinations remain a significant concern [32].
+This implies that artificial intelligence frequently produces inaccurate or deceptive data.
+Many people have doubts regarding the veracity of publications produced by artificial intelligence.
+This hinders the spread of fraudulent or deceptive content and, in many cases, jeopardizes the veracity of information.
+To solve this issue, a multidisciplinary strategy is needed, one that incorporates targeted research to find and fix the root causes of AI system hallucinations.
+If AI-generated material is to become increasingly sophisticated in its ability to distinguish between genuine and fake information, it must pass stringent screening processes and be continuously enhanced.
+In the GenAI age, creating AI content necessitates a constant focus on method improvement and in-depth study to ensure data accuracy.
+A complex network of unanswered concerns is revealed by the ramifications of GenAI technology for ethics, law, and society.
+The proclamation emphasizes how important interdisciplinary collaboration is to this technology’s development and use.
+It entails closely monitoring the ways in which these advancements impact ethical dilemmas, the legal system, and society at general.
+Together, technologists, activists, and the general public must develop a comprehensive plan for the ethical and socially responsible use of artificial intelligence in the digital age.
 6 Discussion
-The sophisticated field of GenAI in cybersecurity has been examined in this paper.
-The focus is on both offensive and defensive strategies.
-By enhancing incident response, automating defensive systems, and identifying sophisticated attacks, GenAI has a disruptive influence that might significantly raise cybersecurity standards.
-Some of the new risks that accompany these technical improvements include hackers having access to ever-more-advanced attack-building tools.
-This discrepancy highlights the significance of striking a balance between purposefully restricting the components that can be employed and enhancing GenAI’s capabilities.
-Apart from the apparent inconsistency between offensive and defensive strategies, this study examines the moral, legal, and societal implications of utilizing artificial intelligence in cybersecurity.
-It also emphasizes the need for robust legal frameworks, strict moral standards, ongoing technical monitoring, and proactive GenAI management.
+This study examined the complex area of GenAI in cybersecurity.
+The two primary areas of emphasis are offensive and defensive strategies.
+By spotting complex assaults, improving incident response, and automating defensive systems, GenAI has the potential to dramatically increase cybersecurity standards.
+These technological advancements give birth to new concerns, such as hackers’ access to ever-more-advanced attack-building tools.
+This contrast highlights how crucial it is to strike a balance between deliberately restricting the components that can be used and enhancing GenAI’s capabilities.
+Moreover, advanced technologies can be combined with GenAI and LLM methods to increase the system’s security posture.
+For example, digital twin technology, which creates digital replicas of physical objects enabling two-way communications [94], can enhance the cybersecurity of systems thanks to its abilities [95].
+This technology can be combined with GenAI methods to boost system resiliency and security.
+In addition to examining the seeming conflict between offensive and defensive strategies, this study looks into the ethical, legal, and social implications of applying AI in cybersecurity.
+It also highlights the necessity of strong moral principles, continuous technical oversight, proactive GenAI management, and strong legal frameworks.
 This is a paradigm-shifting and technical revolution.
-Adopting a holistic strategy considering the technological, ethical, and sociological consequences of implementing GenAI into cybersecurity is crucial.
-Moreover, our findings emphasise the significance of interdisciplinary collaboration to promote GenAI applications in cybersecurity.
+Adopting a holistic strategy that considers the technological, ethical, and sociological consequences of implementing GenAI into cybersecurity is crucial.
+Furthermore, our findings emphasise the significance of interdisciplinary collaboration in promoting GenAI cybersecurity applications.
 The intricacy and findings of GenAI technologies require expertise from various fields, including computer science, law, ethics, and policy-making, to navigate their possible challenges.
@@
 By promoting a consistent focus on the complex relationship between cybersecurity resilience and GenAI innovation, supported by a commitment to ethical integrity and societal advancement, the current study establishes the groundwork for future research initiatives.
+Using innovative technologies and algorithms can help eliminate vulnerabilities in GenAI solutions
 7 Conclusion
-Our thorough examination of cybersecurity offence and defence, as well as Generative Artificial Intelligence (GenAI) technologies, concludes with the discovery of a double-edged sword.
-Although GenAI has the potential to revolutionize cybersecurity processes by automating defences, enhancing threat intelligence, and improving cybersecurity protocols, it also opens up new avenues for highly skilled cyberattacks.
-Incorporating GenAI into cybersecurity emphasises the enduring ethical, legal, and technical scrutiny essential to minimize the risks of misuse and maximize the benefits of this technology for protecting digital infrastructures.
+This work thoroughly examines the Generative Artificial Intelligence (GenAI) technologies in cybersecurity.
+Although GenAI has the potential to revolutionize cybersecurity processes by automating defences, enhancing threat intelligence, and improving cybersecurity protocols, it also opens new vulnerabilities for highly skilled cyberattacks.
+Incorporating GenAI into cybersecurity emphasises the robust ethical, legal, and technical scrutiny essential to minimize the risks of misuse of data and maximize the benefits of this technology for protecting digital infrastructures and systems.
 Future studies should concentrate on creating strong ethical standards and creative defence mechanisms to handle the challenges posed by GenAI and guarantee a fair and impartial approach to its implementation in cybersecurity.
@@
     A.2 Self-replicating simple virus
-    This basic and simple virus can restart the computer (windows as a sample); we didn’t enhance privilege escalation and full antivirus evasion for ethical reasons. See Figure12.
+    This basic and simple virus can restart the computer (windows as a sample); we didn’t enhance privilege escalation and full antivirus evasion for ethical reasons. It can be seen in Fig. 13.
     Figure 13: Self-replicating simple virus
     A.3 Polymorphism
-    This basic and polymorphic design shows that LLMs could assist cyber ops. See Figure13.
+    This basic and polymorphic design shows that LLMs could assist cyber ops.It can be seen in Fig. 14.
     Figure 14: Skeleton code for polymorphic behaviour
@@
     A.4 Rootkit
-    An educational rootkit is developed and improved by GPT3.5 and GPT4. See Figure16.
+    An educational rootkit is developed and improved by GPT3.5 and GPT4. It can be seen in Fig. 17.
     Figure 17: Rootkit
     A.5 Stealthy Data Exfiltration
-    A script for stealthy avoidance of detection by anomaly detection systems was developed and improved by GPT3.5 and GPT4. See Figure17.
+    A script for stealthy avoidance of detection by anomaly detection systems was developed and improved by GPT3.5 and GPT4. It can be seen in Fig. 18.
     Figure 18: Data Exfiltration Script with Stealth Features